HUBBLE SPACE TELESCOPE DAILY REPORT #2143 PERIOD COVERED: 0000Z (UTC) 06/11/98 - 0000Z (UTC) 06/12/98 Daily Status Report as of 163/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Three Sets of NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxies NGC-2989, NGC-4294 and NGC-6744 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.4 Completed Five Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. As described in HSTAR 6525 and in yesterday's report, multiple losses of lock occurred during the third iteration of this proposal. Otherwise, the observations were completed as planned, and no further problems were reported. 1.5 Completed STIS/MA2 7647 (STIS NUV-MAMA Cycle 7 Flats) The Space Telescope Imaging Spectrograph (MA2) was used to make near-ultraviolet (NUV) flat calibration images. This program will obtain NUV-MAMA flat-field observations with the D2 lamp for the construction of on-orbit D-flats for G230M. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of NIC/1/2/3 7871 (The Properties of High Redshift Galaxies : A Near-Infrared Redshift Survey) The Near Infrared Camera (Cameras One, Two and Three) was used to make infrared galactic survey observations of high redshifted galaxies in s sky field designated as Field-1417. A comparison between the population of galaxies in the low {z<1} and high {z>2} redshift universe has shown that the peak era of star forming activity in the universe occurs in the redshift interval 1 1 Cluster Galaxies in the Optical Rest-Frame), wich occurs during the period of the next report, was affected. 2.2 FHST Updates: Scheduled : 13 Successful : 12 The roll delay update at 163/0708Z failed. HSTAR 6531 was written. 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC errors were cleared at 162/1912Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH